OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total 45,292 . The variable costs will be $9 per book. The publisher will sell the finished product to bookstores at a price of $25.75 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales?

OpenStudy (anonymous):

Hi there. Let's start with the easy part. If they sell B books, and make $25.75 per book, what's the total money they take in from book sales?

OpenStudy (anonymous):

(Do you want to work through this to learn it, or just get an answer?)

OpenStudy (lilg132):

45,292 + 9x = 25.75x 45,292 = 25.75x - 9x 45,292 = 16.75x 45,292/16.75 = 16.75x / 16.75 2704 = x so she will need to sell 2704 to break even to check just use the value of x in the equation 45,292 + 9(2704) = 25.75(2704) 45,292 + 24,336 = 69,628 69,628 = 69,628 remember to click good answer

OpenStudy (anonymous):

aww i was doing it backwards!